# Leveraging Photometry for Deconfusion of Directly Imaged Multi-Planet Systems

### v1

@@ Abstract
Planned and future missions, including the Habitable Worlds Observatory (HWO), will aim to directly image Earth-like exoplanets around Sun-like stars in reflected light. Determining whether an exoplanet is in the habitable zone of its star may be difficult in multi-planet systems when the observer does not know in advance which point source detection corresponds to which planet. This “confusion” problem will be a concern for future missions due to the high occurrence rate of multi-planet systems, and will be exacerbated by lack of prior knowledge about planets’ orbital parameters or characteristics, particularly for systems at high inclination with respect to the observer. This work addresses the confusion problem by developing a photometry model and new orbit ranking function to augment the “deconfuser” tool and account for phase variation exhibited by a planet throughout its orbit. We demonstrate the new ranking scheme on a subset of thirty highly confused simulated multi-planet systems among three inclination groupings (low, medium, and high). Results indicate that photometry improves differentiation of previously confused orbits in 7/10 of the low inclination cases, 6/10 of the medium inclination cases, and 6/10 of the high inclination cases. This improvement in handling highly confused systems emphasizes that photometry shows promise for supporting orbit discrimination and deconfusion of directly imaged multi-planet systems, and should be considered when fitting orbits to detections.
Photometry — Orbit determination — Coronagraphic imaging
software: Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), Pandas (The Pandas development team, 2020), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), SciPy (Virtanen et al., 2020)
@@ 1 Introduction
The Astro2020 decadal survey recommended a large IR/O/UV strategic mission to yield direct detections of exoplanets in reflected light at close-in orbital separations (National Academies of Sciences, Engineering, and Medicine, 2021). Such a mission, now deemed the “Habitable Worlds Observatory” (HWO), will directly image exoplanets in reflected light and characterize potentially habitable worlds around Sun-like stars. The goal for such a mission is to reach contrast levels of at least $1\mathrm{e}$ - 1010 (National Academies of Sciences, Engineering, and Medicine, 2021). While the instrument specifications are not yet defined, the inner working angle (IWA) will likely be similar to that of the HabEx Mission Concept (2.4 $\lambda/D$; Gaudi et al. 2020). A coronagraph technology demonstration that will enable further maturation of the technology required to enable HWO will be flown and tested on the Nancy Grace Roman Space Telescope (Roman) (Mennesson et al., 2020). The Roman Coronagraph is expected to achieve contrast ratios of at least $1\mathrm{e}$ - 88 at separations of 3 to 9 $\lambda/D$ (Bailey et al., 2023).
@@
When observing with future direct imaging missions, we may encounter challenges in identifying and characterizing new exoplanets in multi-planet systems. Previous work indicates that a “confusion” problem can arise when multiple detections are made of a multi-planet system and it is not clear which detection belongs to which planet (Keithly & Savransky, 2021; Pogorelyuk et al., 2022). Keithly & Savransky (2021) demonstrated that the Solar System would encounter many incidences of confusion in brightness and separation between planets, especially at higher viewing inclinations. In this work, we refer to “confusion” as the uncertainty in distinguishing between planet detections across multiple epochs of a set of direct imaging observations. The confusion problem can make it difficult to differentiate planets in a system and will exacerbate the orbit-fitting and characterization problem. To address the confusion problem, Pogorelyuk et al. (2022) (hereinafter P22) developed an algorithm called the “deconfuser” [^1] https://github.com/MIT-STARLab/deconfuser which efficiently fits orbits to observations of planets given no prior information on the system.
@@
The deconfuser accepts unlabeled astrometric detections of planets and generates all possible combinations of orbit matches for a system. It specifically provides discrete solutions to match detections to planets in a set of images, within some astrometric error tolerance set by the user. This approach is different than other orbit-fitting tools, such as those that employ Monte Carlo (MC) methods (e.g., Octofitter, Thompson et al. 2023; orbitize!, Blunt et al. 2020, 2024; PyAstrOFit, Wertz et al. 2017) to fit distributions of orbital parameters to planet detections. MC orbit-fitting tools can often be relatively slow due to their iterative processes [^2] For example, testing one set of orbit combinations with orbitize! takes $\sim$6.4 hours on a standard MacBook Pro with 20 temperatures, 500 walkers, and 10,000 steps per walker (see Section 3.2).. Existing orbit-fitting tools, including orbitize!, Octofitter, and PyAstrOFit, also require the user to know which detection belongs to which planet prior to fitting and do not address the detection-confusion problem.
@@
For multi-planet systems, there may be many ways to assign detections to planets, which can lead to numerous combinations of orbits per system that each describe the data equally well. In the deconfuser, orbit matches are initially ranked based on the combinations of orbits that best describe the detections. However, this can still leave many possible combinations of orbits which would benefit from additional ranking steps, particularly for higher inclination systems. Additionally, no assumptions are made about best-fitting orbits in the deconfuser’s initial ranking (e.g., mutual inclination or dynamical stability). After the initial ranking step, it is possible for there to still be multiple assignments of detections that are self-consistent and describe the data equally well.
@@
P22 reported an increased probability ($P$) of confusion for systems with at least three planets and inclinations ($i$) $\geq 75^{\circ}$ when photometry is not taken into account ($P>40\%$ for three observations of a system). Higher confusion rates can result in incorrect orbit assignments which may give very different predictions for multi-planet systems, make planetary characterization difficult, and make observation scheduling more time-consuming. However, highly inclined systems offer the greatest potential for reducing the probability of confusion by accounting for a planet’s brightness variation with phase angle ($\alpha$), due to the ability of the observer to access the full range (for $i=90^{\circ}$) or nearly the full range of orbital phase angles (Bruna et al., 2023). In this work, we build upon the work presented in P22 by developing an additional orbit-ranking step for the deconfuser with the goal of reducing confusion rates by leveraging the phase variation of orbiting planets. We apply the updated orbit ranking scheme on a subset of highly confused simulated systems across low ($i=0-45^{\circ}$), medium ($i=45-70^{\circ}$), and high inclination ($i=70-90^{\circ}$) ranges to demonstrate how photometric information and planetary orbital phase influence the confusion problem across inclination.
@@
Incorporating photometric variation into simulated observations and planet-planet system identification is an involved process; many parameters influence our observations ranging from factors including phase angle variation, individual planet’s atmospheric parameters, instrument systematics, and joint dependence of variables (e.g., astrometry and photometry). Other physical parameters such as rings and moons can also influence exoplanetary spectra and photometry (Arnold & Schneider, 2004; Dyudina et al., 2005; Coulter et al., 2022; Limbach et al., 2024). The scope of this paper is to assume that we have perfect knowledge of planets’ locations in an image and do not include the joint dependence of photometry and relative astrometry. That is, the brightness of a planet detection does not influence our knowledge of the planet’s location in this work. We do not include rings or moons in our simulations. In this work, we assess if and when single-band photometric variation can assist in ranking exoplanet orbit fits using an example subset of highly confused simulated systems.
@@
Accurate orbit determination may also support improved exoplanet characterization. Salvador et al. (2024) showed that prior knowledge of a planet’s orbit and phase angle can improve the constraints on the range of possible radii. Carrión-González et al. (2020) concluded that unknown planetary radius will prevent the detection of clouds and blur other atmospheric information. Improving orbit determination may therefore trickle down to support atmospheric characterization by reducing radius uncertainties.
@@ 1.1 Photometry from Direct Imaging
This section presents a brief overview of exoplanet photometry in the context of our work. Direct imaging provides detections of the light reflected from a planet’s surface and atmosphere, or the self-luminosity of a planet due to internal heat sources, depending on the wavelength of observation. In this work, we assume visible observations of reflected light. Combining reflected light observations with the orbital motion of a planet results in variations in brightness and color due to changing planetary phase angle with respect to the observer (see Figure 1) (e.g., Traub & Oppenheimer 2010; Cahoy et al. 2010; Madhusudhan & Burrows 2012). Previous studies have shown that a planet’s brightness, color, and reflected light spectra change based on the type of planet being observed and the planet’s location in its orbit (e.g., Traub 2003; Mallama 2007, 2009; Cahoy et al. 2010; Madhusudhan & Burrows 2012; Batalha et al. 2018; Mallama & Hilton 2018; Smith et al. 2020). These effects are observed because the properties depend on the geometric albedo and the planet’s phase angle, which are illustrated in Figure 1 and through the following equation for the planet-to-star flux ratio (Sobolev, 1975),
@@
where $\lambda$ is the wavelength of observation. $\alpha$ is the phase angle, where $\alpha=180^{\circ}$ is the angle when the planet is transiting with respect to the observer and $\alpha=0^{\circ}$ is the angle at secondary eclipse (see Figure 1). $A_{g}(\lambda)$ is the wavelength-dependent geometric albedo of the planet at $\alpha=0^{\circ}$ (full-phase), $R_{p}$ is the planet’s radius, $d$ is the planet-star separation, and $\Phi(\lambda,\alpha)$ is the planet’s phase function. The geometric albedo describes the reflectivity of an object at full phase ($\alpha=0^{\circ}$) relative to an identically sized isotropically reflecting surface under the same incident flux. As a planet orbits its host star, the planetary phase change appears to the observer as a change in brightness. In Figure 1, the model phase function is a Lambertian, which assumes diffuse reflection in all directions (see Section 2.1.1; Lambert 1760). We can leverage this photometric variation in the deconfuser to gather information about a planet as it progresses along its orbit, and to assist in breaking ties between orbit options that would otherwise be close to equally ranked.
@@
Figure 1: Planetary orbital phase illustration for an edge-on system ($i=90^{\circ}$) (top), with the phase angle convention used in this work and a simulated Lambertian phase curve (bottom) (see Section 2.1.1 and Equation 3; Brown 2005). Full phase, or $\alpha=0^{\circ}$, occurs behind the star, relative to the observer (secondary eclipse); $\alpha=180^{\circ}$ occurs during the planet’s primary eclipse of the star. $\pm 180^{\circ}$ are the same point on the orbit. We note that the planetary systems explored in this work will not necessarily transit their host star and this figure is for illustrative purposes.
@@
Including atmospheric characteristics or phase variation in the deconfuser could support differentiation of orbit assignments. Previous studies have highlighted the dependence of planetary brightness on orbital separation, location, and atmospheric composition (e.g., Traub 2003; Mallama 2007, 2009; Cahoy et al. 2010; Mallama & Schmude 2012; Gao et al. 2017; Madden & Kaltenegger 2018; Mallama & Hilton 2018; Batalha et al. 2018; Smith et al. 2020). Planetary albedo is dependent on separation due to the amount of flux received and the instellation influencing various condensable species and photochemistry in atmospheres. Previous works have expressed these results through model spectra (e.g., Gao et al. 2017; Hu 2019) and Solar System observations (e.g., Mallama et al. 2006; Mallama & Hilton 2018). This work focuses on the inclusion of brightness information in terms of a planet’s phase variation throughout its orbit. We aim to support the rejection of geometrically valid orbits that are inconsistent with the expectations of planetary appearance.
@@
In this work we model brightness variations throughout an orbit with a single assumed planetary albedo and a Lambertian phase function (see Section 2.1). However, Solar System observations have shown that planetary albedos may vary across epochs (e.g., Lockwood & Jerzykiewicz 2006; Palle et al. 2016; Irwin et al. 2025). Solar System phase functions also deviate from a Lambertian assumption (e.g., Mayorga et al. 2016; Mallama & Hilton 2018; Mayorga et al. 2020). However, Lambertian reflectance generally approximates the shape of a reflected light planetary phase function well, and allows for modeling of the broad shape of a phase curve as a planet orbits its star. We expand on these limitations in Section 4.3 and describe avenues of future work to address these limitations in Section 5.
@@
This work expands on the original version of the deconfuser by developing a new orbit-ranking scheme that uses photometric variation due to orbital phase to help disentangle confused orbit options. We build upon the work in P22 by demonstrating the new photometry ranking scheme on an example subset of highly-confused simulated systems. The confused systems presented in this work are three-planet systems for which the deconfuser is unable to find a single solution to match the detections given only their relative astrometric information. We intentionally target difficult scenarios to demonstrate the photometry ranking scheme, rather than doing a full statistical assessment of potential systems as they could occur randomly in the universe. We demonstrate in Section 3 that single-band photometry shows promise for differentiating orbit combinations for a subset of representative highly confused cases in all but thirty to forty percent of systems across inclination groups.
@@ 1.2 Organization
We outline our approach to including additional ranking metrics for the deconfuser and describe how we test the usefulness of photometry for orbit-fitting of simulated directly imaged multi-planet systems in Section 2. In Section 3, we present the results of a subset of highly confused simulated systems. In Section 3.2, we compare the deconfuser to orbitize! to validate the orbit-fitting methods employed in the deconfuser, and showcase how well it fits orbits to detections compared to the simulated truth. We describe the impact of photometry on deconfusion for these challenging systems and the general performance of the deconfuser in estimating the correct orbital parameters in Section 4. In Section 5, we summarize the results of this work and present directions for future investigations and improvements with the deconfuser.

### v2

@@ Abstract
Planned and future missions, including the Habitable Worlds Observatory, will aim to image Earth-like exoplanets around Sun-like stars in reflected light. Determining whether an exoplanet is in the habitable zone of its star may be difficult in multi-planet systems when the observer does not know in advance which detection corresponds to which planet. This “confusion” problem will be a concern for future missions due to the high occurrence rate of multi-planet systems, and will be exacerbated by lack of prior knowledge about planets’ orbital parameters or characteristics. We address the planet-detection confusion problem by developing a photometry model and new orbit ranking scheme to augment a “deconfuser” tool and account for phase variation exhibited by planets throughout their orbits. We demonstrate the new ranking scheme as a proof-of-concept on a subset of highly confused simulated multi-planet systems among three inclination groupings (low, medium, and high). We find that incorporating photometry improves differentiation of previously confused orbits in more than half of these particularly challenging cases. These results emphasize that photometry is useful for orbit discrimination and deconfusion of directly imaged multi-planet systems, and should be considered when fitting orbits to detections.
keywords:
exoplanets, high-contrast imaging, photometry, data analysis techniques
[…]
@@ 1 Introduction
The Astro2020 decadal survey recommended a large IR/O/UV strategic mission to yield direct detections of exoplanets in reflected light at close-in orbital separations [49]. Such a mission, now deemed the “Habitable Worlds Observatory” (HWO), will directly image exoplanets in reflected light and characterize potentially habitable worlds around Sun-like stars. The goal for such a mission is to reach contrast levels of at least $1\mathrm{e}$ - $10$ [49, 16]. While the instrument specifications are not yet defined, the inner working angle (IWA) will likely be similar to that of the HabEx Mission Concept (2.4 $\lambda/D$ [19]). A coronagraph technology demonstration that will enable further maturation of the technology required to enable HWO will be flown and tested on the Nancy Grace Roman Space Telescope (Roman) [46]. The Roman Coronagraph is expected to achieve contrast ratios of at least $1\mathrm{e}$ - 88 at separations of 3 to 9 $\lambda/D$ [2].
@@
When observing with future direct imaging missions, we may encounter challenges in identifying and characterizing new exoplanets in multi-planet systems. A “confusion” problem can arise when multiple detections are made of a multi-planet system and it is not clear which detection belongs to which planet [31, 53]. Previous work modeling the Solar System demonstrates that our own system’s planets may encounter many incidences of confusion in brightness and separation between planets, especially at higher viewing inclinations [31]. In this work, we refer to “confusion” as the uncertainty in distinguishing between planet detections across multiple epochs of a set of direct imaging observations. The confusion problem can make it difficult to differentiate planets in a system and will exacerbate the orbit-fitting and characterization problem.
@@
To address the confusion problem, Ref. Pogorelyuk et al. [53] developed an algorithm called the “deconfuser” (https://github.com/MIT-STARLab/deconfuser) which efficiently fits orbits to observations of planets given no prior information on the system. The deconfuser was developed to provide a fast orbit-fitting method for flight programs and to address the planet-detection confusion problem. The tool accepts unlabeled astrometric detections of planets and generates all possible combinations of orbit matches for a system. It specifically provides discrete solutions to match detections to planets in a set of images, within some astrometric error tolerance set by the user. This approach is different than other orbit-fitting tools, such as those that employ Monte Carlo (MC) methods (e.g., Octofitter [65], orbitize! [5, 4], PyAstrOFit [70]) to fit distributions of orbital parameters to planet detections. MC orbit-fitting tools also can often be relatively slow due to their iterative processes. For example, testing one set of orbit combinations with orbitize! took us $\sim$6.4 hours on a standard MacBook Pro with 20 temperatures, 500 walkers, and 10,000 steps per walker. For the same system (i.e., three planets, three detections), the deconfuser can process all orbit matches in less than 0.2 seconds [53]. Additionally, existing orbit-fitting tools, including orbitize!, Octofitter, and PyAstrOFit, require the user to know which detection belongs to which planet prior to fitting and do not address the detection-confusion problem. They also do not consider photometric variation due to orbital phase or atmospheric characteristics when returning orbit fits for a given set of detections.
@@
For multi-planet systems, there may be many ways to assign detections to planets, which can lead to numerous combinations of orbits per system that each describe the data equally well. In the deconfuser, orbit matches are initially ranked based on the combinations of orbits that best describe the detections. While relative astrometry is good enough to deconfuse many cases with three to four epochs of observation [53], there are still problematic corner cases where confusion remains and relative astrometric information does not provide enough insight for orbit differentiation. This is true even in the case of perfect knowledge of the detections’ positions (i.e., zero uncertainty). Confusion is a fundamental issue, which can leave many possible combinations of orbits for a set of detections and would benefit from additional ranking steps.
@@
In this work, we build upon the work presented in Ref. Pogorelyuk et al. [53] by developing an additional orbit-ranking step for the deconfuser with the goal of reducing confusion for difficult systems by leveraging the phase variation of orbiting planets. We apply the updated orbit ranking scheme on a set of highly confused simulated systems across low ($i=0-45^{\circ}$), medium ($i=45-70^{\circ}$), and high inclination ($i=70-90^{\circ}$) ranges to assess how we can use photometric information and planetary orbital phase to deconfuse observations of multi-planet systems across inclination ranges. We explore here whether photometry is sufficient for deconfusion in highly confused systems, or if other techniques may be required for orbit differentiation.
@@
Incorporating photometric variation into simulated observations and planet-planet system identification is an involved process; many parameters influence our observations ranging from factors including phase angle variation, individual planet’s atmospheric parameters, instrument systematics, and joint dependence of variables (e.g., astrometry and photometry). Other physical parameters such as rings and moons can also influence exoplanetary spectra and photometry [1, 15, 10, 33]. The scope of this paper is to assume that we have perfect knowledge of planets’ locations in an image and do not include the joint dependence of photometry and relative astrometry. That is, the brightness of a planet detection does not influence our knowledge of the planet’s location in this work. We do not include rings or moons in our simulations. In this work, we assess if and when single-band photometric variation can assist in ranking exoplanet orbit fits using an example subset of highly confused simulated systems. The method presented in this work uses all information that would be immediately available to observers from a set of photometric observations.
@@
Accurate orbit determination may also support improved exoplanet characterization. Previous work showed that prior knowledge of a planet’s orbit and phase angle can improve the constraints on the range of possible radii [57]. Reference Carrión-González et al. [8] concluded that unknown planetary radius will prevent the detection of clouds and blur other atmospheric information. Improving orbit determination may therefore trickle down to support atmospheric characterization by reducing radius uncertainties.
@@ 1.1 Photometry from Direct Imaging
This section presents a brief overview of exoplanet photometry in the context of our work. Direct imaging provides detections of the light reflected from a planet’s surface and atmosphere, or the self-luminosity of a planet due to internal heat sources, depending on the wavelength of observation. In this work, we assume visible observations of reflected light. Combining reflected light observations with the orbital motion of a planet results in variations in brightness and color due to changing planetary phase angle with respect to the observer (see Fig. 1; e.g., Refs. Traub and Oppenheimer [67], Cahoy et al. [7], Madhusudhan and Burrows [36]). Previous studies have shown that a planet’s brightness, color, and reflected light spectra change based on the type of planet being observed and the planet’s location in its orbit (e.g., Refs. Traub [68], Mallama [40, 41], Cahoy et al. [7], Madhusudhan and Burrows [36], Batalha et al. [3], Mallama and Hilton [37], Smith et al. [58]). These effects are observed because the properties depend on the geometric albedo and the planet’s phase angle, which are illustrated in Fig. 1 and through the following equation for the planet-to-star flux ratio [60],
@@
where $\lambda$ is the wavelength of observation. $\alpha$ is the phase angle, where $\alpha=180^{\circ}$ is the angle when the planet is transiting with respect to the observer and $\alpha=0^{\circ}$ is the angle at secondary eclipse (see Fig. 1). $A_{g}(\lambda)$ is the wavelength-dependent geometric albedo of the planet at $\alpha=0^{\circ}$ (full-phase), $R_{p}$ is the planet’s radius, $d$ is the planet-star separation, and $\Phi(\lambda,\alpha)$ is the planet’s phase function. The geometric albedo describes the reflectivity of an object at full phase ($\alpha=0^{\circ}$) relative to an identically sized isotropically reflecting surface under the same incident flux. As a planet orbits its host star, the planetary phase change appears to the observer as a change in brightness. In Fig. 1, the model phase function is a Lambertian, which assumes diffuse reflection in all directions (see Sec. 2.1.1 and Ref. Lambert [32]). We can leverage this photometric variation in the deconfuser to gather information about a planet as it progresses along its orbit, and to assist in breaking ties between orbit options that would otherwise be close to equally ranked.
@@
Figure 1: Planetary orbital phase illustration for an edge-on system ($i=90^{\circ}$) (top), with the phase angle convention used in this work and a simulated Lambertian phase curve (bottom) (see Sec. 2.1.1 and Eq. (3) [6]). Full phase, or $\alpha=0^{\circ}$, occurs behind the star, relative to the observer (secondary eclipse); $\alpha=180^{\circ}$ occurs during the planet’s primary eclipse of the star. $\pm 180^{\circ}$ are the same point on the orbit. We note that the planetary systems explored in this work will not necessarily transit their host star and this figure is for illustrative purposes.
@@
Including atmospheric characteristics or phase variation in the deconfuser could support differentiation of orbit assignments. Previous studies have highlighted the dependence of planetary brightness on orbital separation, location, and atmospheric composition (e.g., Refs. Traub [68], Mallama [40, 41], Cahoy et al. [7], Mallama and Schmude [38], Gao et al. [18], Madden and Kaltenegger [35], Mallama and Hilton [37], Batalha et al. [3], Smith et al. [58]). Planetary albedo is dependent on separation due to the amount of flux received and the instellation influencing various condensable species and photochemistry in atmospheres. Previous works have expressed these results through model spectra (e.g., Refs. Gao et al. [18], Hu [27]) and Solar System observations (e.g., Refs. Mallama et al. [39], Mallama and Hilton [37]). This work focuses on the inclusion of brightness information in terms of a planet’s phase variation throughout its orbit, which is measured along with relative astrometry during photometric observations. We develop an expanded version of the deconfuser with the goal of supporting the rejection of geometrically valid orbits that are inconsistent with the expectations of planetary appearance.
@@
In this work we model brightness variations throughout an orbit with a single assumed planetary albedo and a Lambertian phase function (see Sec. 2.1). However, Solar System observations have shown that planetary albedos may vary across epochs (e.g., Refs. Lockwood and Jerzykiewicz [34], Palle et al. [50], Irwin et al. [29]). Solar System phase functions also deviate from a Lambertian assumption (e.g., Refs. Mayorga et al. [45], Mallama and Hilton [37], Mayorga et al. [44]). Regardless, Lambertian reflectance generally approximates the shape of a reflected light planetary phase function well, and allows for modeling of the broad shape of a phase curve as a planet orbits its star. We expand on these limitations in Sec. 4.3 and describe avenues of future work to address these limitations in Sec. 5.
@@
This work expands the original version of the deconfuser by developing a new orbit-ranking scheme that employs photometric variation due to orbital phase to help disentangle confused orbit options. We build upon the work in Ref. Pogorelyuk et al. [53] by demonstrating the new photometry ranking scheme on an example subset of highly-confused simulated systems. The confused systems presented in this work are three-planet systems for which the deconfuser is unable to find a single solution to match the detections given only their relative astrometric information. We intentionally target difficult scenarios to demonstrate a proof-of-concept with the photometry ranking scheme, since not all system cases will require many orbit-ranking parameters for deconfusion (see confusion rates presented in Ref. Pogorelyuk et al. [53]). We demonstrate in Sec. 3 that single-band photometry shows promise for differentiating orbit combinations in more than half of these representative highly confused cases across inclination groups.
@@ 1.2 Organization
We outline our approach to including additional ranking metrics in the deconfuser and describe how we test whether photometry is sufficient for differentiating orbits fit to simulated directly imaged multi-planet systems in Sec. 2. In Sec. 3, we present the results of the representative highly confused simulated systems. We describe the impact of photometry on deconfusion for these challenging systems and the general performance of the deconfuser in estimating the correct orbital parameters in Sec. 4. In Sec. 5, we summarize the results of this work and present directions for future investigations and improvements with the deconfuser.